OpenStudy (anonymous):

use rational exponents to simplify (^3 have to draw the rest

OpenStudy (anonymous):

\[\sqrt[3]{ab?}^12\]

OpenStudy (anonymous):

that is a 12

OpenStudy (dumbcow):

\[\sqrt[3]{ab}^{12}\]

OpenStudy (anonymous):

tes

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

ok to convert to exponent form....take the index of the radical and make it the denominator \[\sqrt[3]{ab}^{12} = (ab)^{12/3}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you solve that

OpenStudy (dumbcow):

then 12/3 = 4 so (ab)^4

OpenStudy (dumbcow):

there is no solving.....just simplifying

OpenStudy (anonymous):

thats the answer

OpenStudy (dumbcow):

yep